OpenStudy (anonymous):

Find all 3-digit numbers abc (bar on top) with the property that if this number is added to the number obtained by reversing the digits, the sum is 625. ( abc - bar on top- has the digit a in the hundreds place, b in the tens place, and c in the ones place.)

OpenStudy (anonymous):

\[100a+10b+c+100c+10b+a=625\]

OpenStudy (anonymous):

how would i solve that?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

one equation, three variables so infinity many solutions, but they have to be positive integers between 1 and 9 so hmmmm

OpenStudy (anonymous):

we could try some stuff

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

well we could try a = 2, c = 3 so they add up to 5

OpenStudy (anonymous):

then we would have \[200+300+20b+5=625\] and solve for \(b\)

OpenStudy (anonymous):

\[505+20b=625\\ 20b=120 b=6\]

OpenStudy (anonymous):

just grind it till you find it

OpenStudy (anonymous):

typo there

ganeshie8 (ganeshie8):

Ahh that works i guess a=1, c=4 a=2, c=3 a=3, c=2 a=4, c=1 work as well

OpenStudy (anonymous):

yeah you get the same equation for each one \[505+20b=625\]

ganeshie8 (ganeshie8):

abc : 100a + 10b + c cba : 100c + 10b + a ==================== 101(a+c) + 20b = 25*25 no matter what the value of b, 20b always ends up with 0 so a+c needs to be 5 or 15 for the last digit to be 5 on left hand side

OpenStudy (anonymous):

thank you so much!